# Evaluating Prediction-based Interventions

### v1

@@ 2 Related Work
It is well-known, in the experimental economics literature (Banerjee and Duflo, 2009; Deaton and Cartwright, 2018) for example, that insights from RCTs often fail to scale up when the intervention is applied broadly across the entire target population. This failure is often attributed to interaction or interference between units, and has also been referred to, in various contexts, as general equilibrium effects or spillover effects. In the potential outcomes model for causal inference, this is referred to as a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1980, 1990). Recognizing the presence of interference, experimenters typically employ a two-level randomized experiment design (Duflo and Saez, 2003; Crépon et al., 2013) in order to estimate average treatment effects under different treatment proportions (e.g. proportion of job seekers receiving an intervention); treatment proportions are randomly assigned to the treatment clusters, e.g., the labor market. Although there are some exceptions in healthcare (Gohil et al., 2021), such experiment designs have not yet widely been applied to study human-algorithm decision making, where each treatment cluster is associated with a particular human decision-maker.
@@ 3 Existing experimental paradigm and assumptions
In many existing experimental evaluations of algorithmic decision aids, the treatment unit is presumed to be the decision subject and the treatment, $Z_{i}$ is conceptualized as the provision of an predictive risk score to a decision maker who is responsible for making the final decision $D_{i}$. In the case of pre-trial detention decisions, the decision subject is the defendant in a particular court case, the decision maker is the judge presiding on the case, and the treatment $Z_{i}$ is a binary variable indicating if the PSA is shown to the judge for case $i$. $D_{i}$ represents the judge’s binary detention decision, where $D_{i}=1$ means detaining and $D_{i}=0$ means releasing the arrestee before trial. $Y_{i}$ is the binary outcome variable, with $Y_{i}=1$ indicating the arrestee committed an NCA, and $Y_{i}=0$ indicating they did not. $X_{i}$ is a vector of observed pre-treatment covariates for case $i$, including age, gender, race, and prior criminal history. Independence is assumed between each case. We illustrate these variables and their dependencies (or lack thereof) in Figure 1.
@@
The crucial assumption here is that of non-interference between treatment units (see Assumption 1, SUTVA). In words, the treatment status of one unit should not affect the decision for another unit. As a result, $Z_{i}$ is typically randomized at a single-level, i.e., on individual cases; and there is no randomization at the level of decision makers. In Imai et al. (2020) for example, where only one judge participated in the study, the algorithmic score is shown only for cases with an even case number (i.e., $Z_{i}=1$); for odd-numbered cases, no algorithmic recommendation is shown (i.e., $Z_{i}=0$).
The assumption of non-interference across units is questionable, for multiple reasons. For one, in the psychology and behavioural economics literature, it is well-known that decision-makers have a bias towards consistency to their internal policies for decision-making and generally do not change on a case by case basis Tversky and Kahneman (1985). This means, in  Imai et al. (2020) for example, where the judge only sees the risk score for a fraction of the cases, they might choose to ignore the PSA completely on cases where they do see it, or to infer patterns for cases where they don’t see the score, in order to be as consistent as possible in their decision-making (in this work, we mainly model effects such as the former). As such, experiment designs that assume the case-independent model does not allow us to draw clear conclusions on how a judge would modify their default decision-making in response to the algorithmic intervention implemented as a full-scale policy, since we only observe the judge’s decision-making under a partial implementation of the would-be algorithmic intervention.
@@ 4.1 Judge’s decision
Suppose instead that the unit $i$ is treated (i.e., $Z_{i,k}=1$). In this case, the judge receives the algorithmic score $\hat{Y}_{i}$. We denote the recommended decision function as $\bar{D}(y)$ which maps a prediction $y$ to a decision. This function can represent existing decision guidelines used by judges for PSA scores (Stith and Cabranes, 1998).
@@
The judge becomes responsive to the algorithmic recommendation if she encounters the algorithmic recommendation for more units (i.e., greater fraction of units are treated).
@@
$b_{k}$ is the baseline responsiveness and $f\left(\frac{1}{n}\sum_{m=1}^{i}Z_{m,k}\right)$ is the adjustment to the responsiveness based on number of treated units seen so far.
@@
To illustrate, we present a simple thresholding model for $J_{i,k}$ based on the average number of treated units:
@@
5 Results: Characterization of causal experiments under Judge models
@@ 5.1 Violation of SUTVA
Our first observation is that the judge’s changing responsiveness induces interference between units. We will formally show that this results in the violation of SUTVA, stated as follows.
@@
Imai et al. (2020) argued that there was no statistically significant spillover effect in their experiment, hence supporting the assumption of SUTVA. However, the question of whether a conditional randmization test such as a permutation test can detect the violation SUTVA is intricate. Permutation tests typically reorder data to assess the null hypothesis, but in this case, the order does not affect the judge’s average exposure to treatment, $\mathbb{E}[Z_{1,k}]$. Indeed, $J_{i,k}$ is a function of an average over i.i.d. variables as defined in (2), (4), or (5), and it becomes effectively constant for all cases after a certain case count $i>T$, given the law of large numbers. This convergence in $J_{i,k}$ implies that reordering would not alter the joint distribution of $D_{i,k}$’s for $i$ large enough. The conditional randomness test used by Imai et al. (2020), for example, maintains the average treatment proportion, randomizing only the spillover effects that are contingent on case order, and hence cannot effectively detect interference between units that is induced by $J_{i,k}$.
@@ 5.2 Estimation of causal effects
Consider two different treatment assignments (to $n$ total units) that have the same treated and untreated units, but assign these treated units to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). For simplicity, we consider two judges, $k=1,2$, who are assumed to be identical, apart from their assigned cases. Specifically, both of them have decision making model as described in equation (2), where $f(z)=az$ is a linear function and $b_{1}=b_{2}=b$.
@@
Given that units are randomly assigned to treatment (with half of units being treated), $\widehat{ATE}$ appears to be an unbiased estimate of $ATE$. Yet, under our model of $D_{i,k}$ (e.g., (2)), the $ATE$ is clearly under-specified, as $D_{i,k}$ in general depends not only on $Z_{i,k}$ but also $\{Z_{1,k},\cdots,Z_{i-1,k}\}$. If the goal is to estimate the treatment effect of providing a predictive decision aid to a decision maker consistently, we would really like, perhaps, to estimate the average treatment effect under total treatment.
@@ 6 Semi-synthetic Experiments
Using our model of $J_{i,k}$, we can derive an alternate set of judge decisions to that in the original experiment. In the original study, all of the treated and untreated cases were assigned to a single judge. We simulate multiple judges, and a corresponding set of alternative decisions to observe the impact of these biases (Section 4.1) on estimations of the overall average treatment effect estimate on judge decisions. To do this, we change the simulated judge $J_{k}$ assigned to a particular case, and their subsequent decisions $D_{i,k}$, using the models of responsiveness in Equations 2, 4, and 5. Each model is explored in Experiments 1 (Figure 4), 2 (Figure 5), and 3 (Figure 6) respectively. We do not manipulate other case details directly for our simulation, and observe the impact of these changes to the average treatment effect (ATE) of being exposed to a PSA score on decisions. We set $\lambda_{k}(X_{i})$ to be the set of provided original judge decisions, and $\bar{D}(\hat{Y}_{i})$ to be the recommended decision under official interpretation guidelines for the PSA. Descriptive statistics, implementation details and results for decision correctness can be found in Appendix D. The data from  Imai et al. (2020) is publicly available here, and the replication code for experiments is available here. Each result is over a simulation with 1000 trials, and plotted with standard error bars.
Figure 4: Experiment 1: Treatment Exposure Effect. We empirically observe changes to the ATE under different treatment assignments for judges $J_{1}$, $J_{2}$, $J_{3}$. Under this model, if a judge does not respond to the PSA when they see it less than 60% of the time, the ATE is higher for scenarios where at least one judge is guaranteed to see the PSA at or above that frequency. Modifying treatment exposure regimes can adjust the ATE estimate on decisions by up to  10x (for non-white female (NWF) subjects), though this impact varies (i.e. for white males (WM), there is little observed impact).
@@
Figure 5: Experiment 2: Capacity Constraint effect. By modifying the threshold on the PSA score (which ranges from 3 to 6), we can simulate a modification of the decision aid’s positive prediction rate. We find that a lower prediction threshold means that the algorithm predicts detainment at a higher frequency and the measured treatment effect on decisions is lower.
@@
Figure 6: Experiment 3: Low Trust Effect. Under our model, lower accuracy algorithmic recommendations lead to less judge responsiveness and an underestimation of the measured treatment effect on decisions. We model this with hypothetical decisions made under incremental accuracy “boosts” to the algorithmic recommendation.
Conclusion and Future Work
In this work, we describe how experiment design choices can impact human interactions with ADS systems, and importantly, bias our estimates of the treatment effects of algorithmic interventions. Our contribution is primarily theoretical and conceptual, and it remains future work to empirically validate the models of human decision making that we have proposed. We hope this work will spark a conversation about choices and challenges in experimental design in the context of algorithmic interventions, with the goal of better measuring and addressing the societal impact of such deployments in consequential real-world settings.
@@ Appendix D Experiment Details
In order to investigate this, we conduct an empirical study similar to that described in Section LABEL:. Namely, for each study, we consider two different treatment assignments (to $n$ total units) that have the same treated and untreated units, but assign these treated units to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). We consider $k$ judges, who are assumed to be identical, apart from their assigned cases. Specifically, all judges have the decision making model as described in equation (2), though we also explore additional experiments where $f(z)=az$ is a simple linear function and $b_{1}=b_{2}=b$.

### v2

@@ 2 RELATED WORK
It is well-known, in the experimental economics literature (Banerjee and Duflo, 2009; Deaton and Cartwright, 2018) for example, that insights from RCTs often fail to scale up when the intervention is applied broadly across the entire target population. This failure is often attributed to interaction or interference between decision subjects, and has also been referred to, in various contexts, as general equilibrium effects or spillover effects. In the potential outcomes model for causal inference, this is referred to as a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1980, 1990). Recognizing the presence of interference, experimenters typically employ a two-level randomized experiment design (Duflo and Saez, 2003; Crépon et al., 2013) in order to estimate average treatment effects under different treatment proportions (e.g. proportion of job seekers receiving an intervention); treatment proportions are randomly assigned to the treatment clusters, e.g., the labor market. Although there are some exceptions in healthcare (Gohil et al., 2021), such experiment designs have not yet widely been applied to study human-algorithm decision making, where each treatment cluster is associated with a particular human decision-maker.
@@ 3 EXISTING EXPERIMENTAL PARADIGM AND ASSUMPTIONS
In many existing experimental evaluations of algorithmic decision aids, the treatment unit is presumed to be the decision subject and the treatment, $Z_{i}$ is conceptualized as the provision of an predictive risk score $\hat{Y}_{i}$ to a decision maker who is responsible for making the final decision $D_{i}$. In the case of pre-trial detention decisions, the decision subject is the defendant in a particular court case, the decision maker is the judge presiding on the case, and the treatment $Z_{i}$ is a binary variable indicating if the PSA is shown to the judge for case $i$. $D_{i}$ represents the judge’s binary detention decision, where $D_{i}=1$ means detaining and $D_{i}=0$ means releasing the arrestee before trial. $Y_{i}$ is the binary outcome variable, with $Y_{i}=1$ indicating the arrestee committed an NCA, and $Y_{i}=0$ indicating they did not. $X_{i}$ is a vector of observed pre-treatment covariates for case $i$, including age, gender, race, and prior criminal history. Independence is assumed between each case. We illustrate these variables and their dependencies (or lack thereof) in Figure 1.
@@
The crucial assumption here is that of non-interference between decision subjects (see Assumption 1, SUTVA). In words, the treatment status of one unit should not affect the decision for another unit. As a result, $Z_{i}$ is typically randomized at a single-level, i.e., on individual cases; and there is no randomization at the level of decision makers. In Imai et al. (2020) for example, where only one judge participated in the study, the algorithmic score is shown only for cases with an even case number (i.e., $Z_{i}=1$); for odd-numbered cases, no algorithmic recommendation is shown (i.e., $Z_{i}=0$).
The assumption of non-interference across subjects is questionable, for multiple reasons. For one, in the psychology and behavioural economics literature, it is well-known that decision-makers have a bias towards consistency to their internal policies for decision-making and generally do not change on a case by case basis Tversky and Kahneman (1985). This means, in  Imai et al. (2020) for example, where the judge only sees the risk score for a fraction of the cases, they might choose to ignore the PSA completely on cases where they do see it, or to infer patterns for cases where they don’t see the score, in order to be as consistent as possible in their decision-making (in this work, we mainly model effects such as the former). As such, experiment designs that assume the case-independent model does not allow us to draw clear conclusions on how a judge would modify their default decision-making in response to the algorithmic intervention implemented as a full-scale policy, since we only observe the judge’s decision-making under a partial implementation of the would-be algorithmic intervention.
@@ 4.1 Judge’s decision
Suppose instead that the unit $i$ is treated (i.e., $Z_{i,k}=1$). In this case, the judge receives the algorithmic score $\hat{Y}_{i}$. We denote the recommended decision function as $\bar{D}(y)$ which maps a prediction $y$ to a decision. This function can represent existing decision guidelines used by judges for PSA scores (Stith and Cabranes, 1998). Note that prediction (lower-case) $y$ in $\bar{D}(y)$ is a variable that can take on any value $\hat{Y}_{i}$. This is to distinguish $\bar{D}(y)$ (the recommended decision, which is a function of the prediction), and $D(z)$ (the actual decision, which is a function of the judge being treated or not).
@@
The judge becomes responsive to the algorithmic recommendation if she encounters the algorithmic recommendation for more subjects (i.e., greater fraction of subjects are treated).
@@
$b_{k}$ is the baseline responsiveness and $f\left(\frac{1}{n}\sum_{m=1}^{i}Z_{m,k}\right)$ is the adjustment to the responsiveness based on number of treated decision subjects seen so far.
@@
To illustrate, we present a simple thresholding model for $J_{i,k}$ based on the average number of treated decision subjects:
@@
5 RESULTS
@@ 5.1 Violation of SUTVA
Our first observation is that the judge’s changing responsiveness induces interference between decision subjects. We will formally show that this results in the violation of SUTVA, stated as follows.
@@
Imai et al. (2020) argued that there was no statistically significant spillover effect in their experiment, hence supporting the assumption of SUTVA. However, the question of whether a conditional randmization test such as a permutation test can detect the violation SUTVA is intricate. Permutation tests typically reorder data to assess the null hypothesis, but in this case, the order does not affect the judge’s average exposure to treatment, $\mathbb{E}[Z_{1,k}]$. Indeed, $J_{i,k}$ is a function of an average over i.i.d. variables as defined in (2), (4), or (5), and it becomes effectively constant for all cases after a certain case count $i>T$, given the law of large numbers. This convergence in $J_{i,k}$ implies that reordering would not alter the joint distribution of $D_{i,k}$’s for $i$ large enough. The conditional randomness test used by Imai et al. (2020), for example, maintains the average treatment proportion, randomizing only the spillover effects that are contingent on case order, and hence cannot effectively detect interference between decision subjects that is induced by $J_{i,k}$.
@@ 5.2 Estimation of causal effects
Consider two different treatment assignments (to $n$ total decision subjects) that have the same treated and untreated decision subjects, but assign these treated decision subjects to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). For simplicity, we consider two judges, $k=1,2$, who are assumed to be identical, apart from their assigned cases. Specifically, both of them have decision making model as described in equation (2), where $f(z)=az$ is a linear function and $b_{1}=b_{2}=b$.
@@
Given that decision subjects are randomly assigned to treatment (with half of the decision subjects being treated), $\widehat{ATE}$ appears to be an unbiased estimate of $ATE$. Yet, under our model of $D_{i,k}$ (e.g., (2)), the $ATE$ is clearly under-specified, as $D_{i,k}$ in general depends not only on $Z_{i,k}$ but also $\{Z_{1,k},\cdots,Z_{i-1,k}\}$. If the goal is to estimate the treatment effect of providing a predictive decision aid to a decision maker consistently, we would really like, perhaps, to estimate the average treatment effect under total treatment.
@@ 6 SEMI-SYNTHETIC EXPERIMENTS
Using our model of $J_{i,k}$, we can derive an alternate set of judge decisions to that in the original experiment. In the original study, all of the treated and untreated cases were assigned to a single judge. We simulate multiple judges, and a corresponding set of alternative decisions to observe the impact of these biases (Section 4.1) on estimations of the overall average treatment effect estimate on judge decisions. To do this, we change the simulated judge $J_{k}$ assigned to a particular case, and their subsequent decisions $D_{i,k}$, using the models of responsiveness in Equations 2, 4, and 5. Each model is explored in Experiments 1 (Figure 4), 2 (Figure 5), and 3 (Figure 6) respectively. We do not manipulate other case details directly for our simulation, and observe the impact of these changes to the average treatment effect (ATE) of being exposed to a PSA score on decisions. We use a binary normalized form of judge decisions (i.e. $D_{i,k}=0$ for a signature bond release, $D_{i,k}=1$ for any kind of cash bond). We set $\lambda_{k}(X_{i})$ to be the set of provided original judge decisions, and $\bar{D}(\hat{Y}_{i})$ to be the recommended decision under official interpretation guidelines for the PSA. Descriptive statistics, implementation details and results for decision correctness can be found in Appendix D. The data from  Imai et al. (2020) is publicly available here, and the replication code for experiments is available here. Each result is over a simulation with 1000 trials, and plotted with standard error bars.
Note that the average treatment effects reported for the empirical results are not absolute, but are changes in the measured treatment effect relative to the baseline reported treatment effect in the original single-judge study. As such, we already have a baseline as defined by the original experiment and specifically report deviations from this baseline as our main results.
6.1 Experiment 1: Treatment Exposure Effect
As discussed in Appendix  D, we investigate two cases – one with two judges (treated at 50% each (Scenario 1A) or  100% treated and  0% treated (Scenario 1B)), and one with three judges (treated at 33% each (Scenario 2A) or at 60%, 30% and 10% respectively (Scenario 2B)). We observe that the largest treatment effect measurements on decisions and decision correctness (ie. the correlation of the decision and the recorded downstream outcome) apply in Scenario 1B, though differences are smaller in the latter case. We find that modifying treatment exposure regimes can adjust the causal effect estimate on decisions by up to ${}10x$ (i.e. for non-white female (NWF) subjects), though this impact varies (e.g. for white men (WM), there is little observed impact).
Figure 4: Experiment 1: Treatment Exposure Effect. We empirically observe changes to the ATE under different treatment assignments for judges $J_{1}$, $J_{2}$, $J_{3}$. Results for 1000 trials, with M = Male, F = Female, NW = Non-White, W = White.
6.2 Experiment 2: Capacity Constraint effect
The available data allows decision thresholds to be set at a level of 3,4,5 or 6. Given the same treatment exposure, we find that there are average treatment effect measurement differences under different algorithm decision thresholds (and thus changes to the positive prediction rate of the algorithm). For example, at a lower algorithmic decision threshold (eg. threshold = 3), the algorithm predicts $\hat{Y}=1$ (i.e. detainment) at a higher frequency and, under our judge responsiveness model, the measured treatment effect on decisions is lower. Interestingly, this effect does not meaningful impact the measure of changes to simulated decision correctness (See details in Appendix  D).
@@ 6.2 Experiment 2: Capacity Constraint effect
Figure 5: Experiment 2: Capacity Constraint effect. By modifying the threshold on the PSA score (which ranges from 3 to 6), we can simulate a modification of the decision aid’s positive prediction rate. We find that a lower prediction threshold means that the algorithm predicts detainment at a higher frequency and the measured treatment effect on decisions is lower. Results are for 1000 trials, with M = Male, F = Female, NW = Non-White, W = White. In this illustrative example, $J_{1}$, $J_{2}$ and $J_{3}$ are assigned $Z=1$ in about 33% of their cases respectfully.
6.3 Experiment 3: Low Trust Effect
In the low trust setting, we derive hypothetical decisions made under different algorithmic recommendation accuracy “boosts”, where an accuracy boost of 0 results in the default algorithmic recommendation accuracy of the real world setting, with respect to actual real world outcomes ($\approx 54\%$). If the boost is at 0.1, then 10% of the algorithmic recommendations are flipped to the actual correct decision with respect to the recorded real world outcome - this simulates a higher accuracy model (eg. the 10% boosted algorithmic recommendation accuracy is $\approx 67\%$). Under our judge responsiveness model in this setting, low accuracy predictions lead to an underestimation of the measured treatment effect on decisions and this impact is even greater for decision correctness (See details in Appendix  D).
@@ 6.3 Experiment 3: Low Trust Effect
Figure 6: Experiment 3: Low Trust Effect. Under our model, lower accuracy algorithmic recommendations lead to less judge responsiveness and an underestimation of the measured treatment effect on decisions. We model this with hypothetical decisions made under incremental accuracy “boosts” to the algorithmic recommendation. Results are for 1000 trials, with M = Male, F = Female, NW = Non-White, W = White. In this illustrative example, $J_{1}$, $J_{2}$ and $J_{3}$ are assigned $Z=1$ in about 33% of their cases respectfully.
7 CONCLUSION
Past work looking at judge responsiveness to automated decision systems explicitly models the phenomenon as a fixed, static property of the judge interacting with the prediction used for automated decision support. This applies to past work in HCI (eg,  Green and Chen (2019c) assume a fixed judge bias, similar to $b_{k}$) and past work on observational studies done in economics (eg.  Albright (2019); Hoffman et al. (2018b) and others also assume a fixed judge bias, similar to $b_{k}$). In this work, we propose that the variability in judge responsiveness is not just a fixed attribute inherent to each judge, but that this judge variability likely also arises as the result of the impact of specific experiment design decisions on overall judge responsiveness (i.e. experiment design choices that impact treatment exposure, model trust, and capacity constraints); judge responsiveness is thus not just a static property inherent to each judge, but also impacted by conditions that the experimenter can control. We thus integrate our understanding of interference in causal inference with the existing discussion on judge responsiveness, demonstrating that if we model the prediction as part of an intervention, estimates of causal effects will be influenced by experiment design decisions through the mechanism of the judge’s decision-making across cases.
This work has implications on how we design future experiments in the context of prediction-based interventions, as well as how we study and measure treatment effects in the context of experimental evaluations for prediction-based interventions. Further study of multi-judge experimental settings, two-level randomization designs, and multi-factor experiment designs (i.e., allowing for comparisons of multiple prediction system options, multiple thresholds, and thus positive prediction rates, etc.) for prediction-based interventions is needed. Further theoretical directions to explore include further discussions on the interactions between multiple responsiveness factors on overall treatment effects, long-term learning effects, group-level impacts, as well as the exploration of measurement corrections to the reported deviations to causal effect estimates. We truly hope that many of these directions can be explored further in future work.
Although our empirical results are limited to data from the judicial pre-trial risk assessment context, our justification for the judge responsiveness factors derives from empirical and theoretical evidence observed across a variety of domains, including healthcare and education (see Appendix  E.2,  E.1). The judge responsiveness factors thus apply more broadly than the criminal justice domain and have been previously observed in other settings. We hope to see further exploration of these other domains in future work.
Acknowledgments
I.D.R. was funded by the Mozilla Foundation and the MacArthur Foundation. L.T.L. was supported by an Amazon Research Award Fall 2023. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of Amazon. We appreciate the comments and feedback from colleagues Avi Feller, Benjamin Recht and Imai Kosuke.
@@ Appendix D Experiment Details
In order to investigate this, we conduct an empirical study similar to that described in Section 5.2. Namely, for each study, we consider two different treatment assignments (to $n$ total decision subjects) that have the same treated and untreated decision subjects, but assign these treated decision subjects to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). We consider $k$ judges, who are assumed to be identical, apart from their assigned cases. Specifically, all judges have the decision making model as described in equation (2), though we also explore additional experiments where $f(z)=az$ is a simple linear function and $b_{1}=b_{2}=b$.
@@ D.1 Simulation Details
Note that the measured outcome of our simulations is not the absolute average treatment effect across all judges, but the change in average treatment effect from the baseline effect measured in the decisions from the original experiment – all indications in figures or text indicated "Estimated Average Effects" are thus indicative of "Estimated Deviation of Average Effects" from the reported baseline.
@@
Clarifications on Scope
Our empirical results only focus on data from the pre-trial risk assessment context for various reasons: as the authors state in their paper,  Imai et al. [2020] is likely “to the best of our knowledge, this is the first real-world randomized controlled trial (RCT) that evaluates the impacts of algorithmic risk assessment scores on judges’ decisions in the criminal justice system”. The field of experimental evaluation for algorithmic deployments is increasing but still nascent – many studies are observational or rely on a benchmark framework. As such, this is one of very few academic experimental evaluations of an ADS system and the only one we could identify for which the full experiment results are publicly available. As we discuss in Appendix  E.2 on Additional References, there are experimental evaluations done in other fields (notably education, and healthcare), but none of these existing evaluations have a tractable release of data. For healthcare, for example, clinical trial data needs to be registered and requested through official government channels, and approval is needed for additional synthetic manipulation. That being said, in future work, through the continued explorations of further partnerships and collaborations, we hope to be able to obtain data from such settings, where there are multiple judges (ie. doctors, teachers, hiring managers) available in the data. This would allow us to further empirically validate some of our assumptions on judge responsiveness. Also, as we mention in Appendix  E.1, our justification for the judge responsiveness factors derives from empirical and theoretical evidence referenced to be observed across a variety of domains – thus, we know from these historical and academic observations that the judge responsiveness factors apply more broadly than the criminal justice domain, and has been previously observed in other settings.